B 25 35 9 R P A 12 C

what are my i known side lengths with detailed explanation
Answer #1

Two triangles are said to be similar if their corresponding angles are congruent and the corresponding sides are in proportion . In other words, similar triangles are the same shape, but not necessarily the same size. The triangles are congruent if, in addition to this, their corresponding sides are of equal length.

Therefore from figure,

In ∆QPR and ∆BAC,

QP/BA = PR/AC = QR/BC

According to the question,

a/9 = b/12 = 25/15

a = (25/15) *(9)

a = 15

b = (25/15)*(12)

b = 20

Please rate my answer with thumbs up if it helps!